OpenStudy (anonymous):

log (x+2) base 2-log (x-5) base 2=3

OpenStudy (amoodarya):

\[\log_{b} x-\log_{b} y=\log_{b} \frac{ x }{ y }\] apply it now then simplify

OpenStudy (amoodarya):

\[\log_{2}(x+2)-\log_{2}(x-5)=3\\log_{2}\frac{ x+2 }{ x-5 }=3 \log_{2}2\\log_{2}\frac{ x+2 }{ x-5 }= \log_{2}2^3\\\log_{2}\frac{ x+2 }{ x-5 }= \log_{2}8\\ \frac{ x+2 }{ x-5 }=8\]